OpenStudy (anonymous):

can someone help me check my work for this problem: y^2(y^2 -4) = x^2(x^2 -5) ...implicitly differentiate and then find equation of tangent line at point (0, -2)

OpenStudy (anonymous):

when i worked it out i got y'=(8x^3 -10x)/(8y^3 - 8y)

OpenStudy (anonymous):

then i plugged in x for all x values and y for all y values to get the value for slope to use the point-slope formula....i ended up getting y = 2 because the numerator was = 0

OpenStudy (turingtest):

\[y^4-4y^2=x^4-5x^2\] take the derivative w/respect to x: \[4y^3y'-8yy'=2x^3-10x=y'(4y^3-8y)\] \[y'=(2x^3-10x)/(4y^3-8y)\]

OpenStudy (turingtest):

plug in the values

OpenStudy (anonymous):

hmm...i wonder how i got such different values...when i first started i distributed the y^2 and the x^2 to the terms in the parenthesis..would that make a difference?

OpenStudy (anonymous):

i ended up taking derivative of: 2y^4 - 4y^2 for the y^2(y^2 -4) part

OpenStudy (anonymous):

after i distributed that y^2 to those terms, i took der. by the sum rule

OpenStudy (turingtest):

That was my first step too, but I don't see how you got 2y^4; where did the 2 coefficient come from? As you can see though at (0,-2) we still get y'=0 y+2=0(x-0) y=-2 so your answer seems right to me anyway.

OpenStudy (anonymous):

ahhhh i distributed it incorrectly haha! thanks for the help im glad i asked to have someone review it haha, thanks again!

OpenStudy (turingtest):

oh yeah, and you had y=2 when it should be y=-2 Anytime!

OpenStudy (anonymous):

thanks =D